# The Hasse Principle for Geometric Variational Problems: An Illustration via Area-minimizing Submanifolds

### v4

@@ Abstract.
The Hasse principle in number theory states that information about integral solutions to Diophantine equations can be pieced together from real solutions and solutions modulo prime powers. We show that the Hasse principle holds for area-minimizing submanifolds: information about area-minimizing submanifolds in integral homology can be fully recovered from those in real homology and mod n homology for all $n\in\mathbb{Z}_{\geq 2}$. As a consequence we derive several surprising conclusions, including: area-minimizing submanifolds in mod n homology are asymptotically much smoother than expected, area-minimizing submanifolds are not generically calibrated, and products of area-minimizing submanifolds are not generically area-minimizing. We conjecture that the Hasse principle holds for all geometric variational problems that can be formulated on chain space over different coeffiicients, e.g., Almgren-Pitts min-max, mean curvature flow, Song’s spherical Plateau problem, minimizers of elliptic and other general functionals, etc.
@@ 1. Introduction
The Hasse principle [19, 44] in number theory states that information about integral solutions to Diophantine equations can be pieced together from real solutions and all solutions modulo prime powers. An illustrative example is the Hasse-Minkowski theorem for quadratic forms [43]. Though the Hasse principle does not hold in general [42], the guiding philosophy of “local-to-global” reasoning remains a cornerstone of modern number theory. Prominent examples include the original formulation of the Birch and Swinnerton-Dyer conjecture [3], where the number of solutions modulo primes conjecturally gives ranks of rational solutions of elliptic curves, and Siegel’s weight formula [12], where the number of solutions modulo prime powers yield weighted sums of number of solutions over $\mathbb{Z}$.
@@
When the author was an undergraduate at Duke University, Professor William Allard recalled a conversation he once had with Professor Herbert Federer, in which they discussed whether the title of the fourth manuscript, Homological Integration Theory, of the monumental work [13], might be too grand a name for its actual contents. Federer remarked that, in his view, homology constituted the most essential component of the entire theory.
@@ Acknowledgements
When the author was an undergraduate at Duke, Professor Allard told many stories of the early days of geometric measure theory. One night, the author had a dream, where Professor Almgren was discussing flat chains with Professor Allard, with Professor Almgren emphasizing that “the key is to change the coefficients.”  That dream sparked a long-standing quest to find the meaning of the enigmatic quote, a pursuit that ultimately culminated in this work.
The author is deeply indebted to Professor Simon for his extraordinary generosity, guidance, and kindness throughout the years. On the occasion of his 80th birthday, the author warmly extends best wishes to Professor Simon. Sincere thanks are paid to Professors Frank Morgan and Antoine Song, whose works [47, 32] are the inspirations for the author to go from local regularity to global geometry. Professor Morgan has also raised several suggestions that ultimately inspire Section 7.1. The author would also like to thank Professor Chao Li, whose questions convince the author to write this manuscript. Special thanks go to Professor Yang Li, who have raised many helpful suggestions. He pointed out the example of Siegel mass formula and suggested the author to keep the discussion around Theorem 4.

### v5

@@ Abstract.
The Hasse principle in number theory states that information about integral solutions to Diophantine equations can be pieced together from real solutions and solutions modulo prime powers. We show that the Hasse principle holds for area-minimizing submanifolds: information about area-minimizing submanifolds in integral homology can be fully recovered from those in real homology and mod $n$ homology for all $n\in\mathbb{Z}_{\geq 2}$. As a consequence we derive several surprising conclusions, including: area-minimizing submanifolds in mod $n$ homology are asymptotically much smoother than expected, area-minimizing submanifolds are not generically calibrated, and products of area-minimizing submanifolds are not generically area-minimizing. We conjecture that the Hasse principle holds for all geometric variational problems that can be formulated on chain spaces over different coeffiicients, e.g., Almgren-Pitts min-max, mean curvature flow, Song’s spherical Plateau problem, minimizers of elliptic and other general functionals, etc.
@@ 1. Introduction
The Hasse principle [19, 44] in number theory states that information about integral solutions to Diophantine equations can be pieced together from real solutions and all solutions modulo prime powers. An illustrative example is the Hasse-Minkowski theorem for quadratic forms [43], and its vast generalization Siegel’s weight formula [12], the latter which of uses the number of solutions modulo prime powers to give a weighted sums of number of solutions over $\mathbb{Z}$. Another example is the original formulation of the Birch and Swinnerton-Dyer conjecture [3], where the number of solutions modulo primes conjecturally gives ranks of rational solutions of elliptic curves. Though the Hasse principle does not hold in general [42], the guiding philosophy of “local-to-global” reasoning remains a cornerstone of modern number theory.
@@
When the author was an undergraduate at Duke University, Professor William Allard recalled a conversation he once had with Professor Herbert Federer in which they discussed whether the title, Homological Integration Theory, of the fourth chapter in the draft of the monumental work [13], might be too grand a name for its actual contents. Federer remarked that, in his view, homology constituted the most essential component of the entire theory.
@@ Acknowledgements
The author would like thank his advisor Professor Camillo De Lellis for his unwavering support. Many helpful suggestions about this manuscript were due to him. Special tribute is paid to Professor William Allard and Professor Frederick Almgren. When the author was an undergraduate at Duke, Allard told many stories of the early days of geometric measure theory, including his conversation with Federer. One night the author had a dream, where Almgren was discussing flat chains with Allard, with Almgren emphasizing that “the key is to change the coefficients.” That dream sparked a long-standing quest to find the meaning of the enigmatic quote, a pursuit that ultimately culminated in this manuscript.
The author is deeply indebted to Professor Leon Simon for his extraordinary generosity, guidance, and kindness throughout the years. On the occasion of his 80th birthday, the author warmly extends best wishes to Professor Simon. Sincere thanks are paid to Professors Frank Morgan and Antoine Song, whose works [47, 32] are the inspirations for the author to go from local regularity to global geometry. Professor Morgan has also raised several suggestions that ultimately inspire Section 7.1. A conversation with Professor Michael Freedman partially inspired Theorem 3. The author would also like to thank Professor Chao Li, whose questions convince the author to write this manuscript. Special thanks go to Professor Yang Li, who have raised many helpful suggestions. He pointed out the example of Siegel mass formula and suggested the author to write the discussion around Theorem 4.